# Mind the Prompt: A Novel Benchmark for Prompt-based Class-Agnostic Counting

### v1

@@ 3.3 Mosaic test
Differently from the negative test, in the mosaic test, counting models are asked to count objects in a query image where two classes are present. The objective is, therefore, to quantitatively assess the ability of the model to distinguish between the correct class and another randomly chosen class that is present in the same image. This test wants to emulate real case scenarios in which an image contains multiple object classes, and the model is requested to count only one. Unfortunately, suitable real-world counting datasets annotated with multiple classes are difficult to obtain. Therefore, from largely used datasets like FSC-147 containing images with a single class of objects, we artificially create multi-class images by concatenating a positive class image with a negative class one, similarly to [2].
@@
Specifically, we are given all the possible pairs of images from the dataset $\{I_{i},I_{j}\}_{\begin{subarray}{c}i,j=1\\ j\neq i\end{subarray}}^{N}$, and the positive class $P_{i}$ of the image $I_{i}$. We call $I_{i}$ and $I_{j}$ the positive and the negative images, respectively. We create the mosaics $I_{ij}^{\text{coll}}=\text{vstack}(I_{i},I_{j})$, which are obtained by stacking the positive image on top of the negative ones on the $y$ axis. We then perform inference with the target counting model using the positive label $P_{i}$ – associated with the top image in the collage – in the following way: $c_{ij}^{\text{pos}},c_{ij}^{\text{neg}}=\mathcal{M}(I_{ij}^{\text{coll}},P_{i})$. Specifically, the returned $c_{ij}^{\text{pos}}$ is the count relative to the top part of the collage, while $c_{ij}^{\text{neg}}$ is the count relative to the below image of the collage. This can be easily obtained from most density-based models, as it is sufficient to cut the output density map in half along the y-axis and integrate them separately to obtain the respective positive and negative counts. Notice that, in this test, the positive count $c_{ij}^{\text{pos}}$ depends on $j$ as well (the different negatives), as the model may be distracted to count the positive instances depending on which negative image is present in the bottom part of the mosaic. Similarly to the negative test, we would require $c_{ij}^{\text{pos}}$ to converge to the ground truth count, while $c_{ij}^{\text{neg}}$ to converge identically to zero, as the model (i) should not be fooled by the negative examples while attending the positive instances and (ii) should not count any of the negative instances in the collage. We summarize this inference procedure in the right part of Figure 2.
@@
The metrics for assessing such behavior should evidence whether the model correctly counts only the correct class while ignoring elements from the negative one. Due to the presence of positive and negative instances for each probed mosaic, this task has many similarities with object detection, where we may have correct and incorrect detections. In this scenario, usually, the performance is computed by using precision – which accounts for the number of correct predictions among all the model outputs – and recall – which represents the fraction of correct instances that were retrieved among all the ground-truth correct instances. Differently from the detection case, in our scenario, we do not have precise information about the correctness of each proposed instance, given that we do not know – and we are not requested to know – the precise localization of each instance within the image, as we are only interested in the final aggregated count for the provided class of objects. Such information could be partially recovered from the two output counts $c_{ij}^{\text{pos}}$ and $c_{ij}^{\text{neg}}$ by making some reasonable assumptions that we will present in the following paragraphs, where we present the precision-recall adapted metrics for the counting task.
Counting Precision and Recall.
@@
Under the above considerations and assumptions, we obtain that counting precision (CntP) and counting recall (CntR), for each mosaic image, can be defined as follows:
@@
We refer the reader to the supplementary material for the derivation of such expressions from the idea shown in Figure 3.
@@
Counting Balanced F-score.
A common way to aggregate precision and recall to obtain a single indicator is the F-score, which is defined as the harmonic mean of the two metrics. We therefore derive a counting F1-score (CntF1) indicator, as follows:
@@ 4.1 Dataset and Methods
Our benchmark is based on the widely-used FSC-147 dataset, designed for class-agnostic object counting that includes a total of 6,135 images. The training portion of FSC-147 consists of images from 89 distinct categories, while both the validation and test sets are composed of images from 29 different categories. Since, in our work, we probe pre-trained state-of-the-art models, we only consider the test and validation splits composed of 1190 and 1286 images, respectively. Notably, there is no overlap between the classes in the training, validation, and test sets, positioning FSC-147 as an open-world dataset. Although most of the images in FSC-147 contain only one class of objects, there are a few containing multiple classes. In order to avoid interfering with our negative-examples evaluation, we filter all the images from test and validation sets where there are more than two object classes. In this way, we are sure that there are no false positive contributions from the other classes in the same image.
@@
We employed the code and the pre-trained models provided by the authors. We stuck to the original image preprocessing pipelines, and we kept the author’s chosen hyperparameters and prompting schema. We only needed to adjust the DAVE inference procedure to behave correctly on our benchmark. In the supplementary material, we better comment on this small yet important modification, also reporting the PrACo performance of the original DAVE implementation.
@@
Figure 5: This figure shows, for each model, the output density maps for three different (mosaic, input prompt) pairs. In each figure, the count reported in the blue box is $c^{\text{pos}}_{ij}$, while the count reported in the red box corresponds to $c^{\text{neg}}_{ij}$. We can notice how the models incorrectly attend to the instances from the negative image in the mosaic, although most of them provide a good estimate for the positive instances in the upper image of the mosaic.
@@ 4.2 Results
We report the results for both the test and validation splits in Tables 1 and 2, respectively. As we can notice, although the methods achieve remarkable performance on standard counting error measures (MAE and RMSE), the outcome is quite heterogeneous on the PrACo metrics. Notably, the negative test on one-stage methods like CounTX, CLIP-Count, and VLCounter shows that the average negative count is comparable to the ground-truth count of the correct class ($\text{NMN}\approx 1$), with CLIP-Count and VLCount achieving $\text{NMN}>1$. This trend is further confirmed by the PCCN metric, where CLIP-Count has the correct estimate nearer to the ground-truth only 38% of the times on the test set. Instead, the two-stage detectors TFPOC and DAVE achieve the best performance on the negative test. The same trends are mostly confirmed on the mosaic test. Specifically, while the CntR metric only increments by around 12% between the worst-performing method (CounTX) and the best-performing one (DAVE), the CntP shows a notable increase, from 0.517 of VLCounter to 0.843 of DAVE, which is an improvement of around 63%. This shows how the methods, despite being mostly able to count the correct class, have a highly heterogeneous behavior with respect to the negative images within the mosaics. It also worth noticing that more recent methods do not always improve on the PrACo metrics, meaning that this aspect is still largely underestimated. At the same time, it is interesting to observe how TFPOC, a training-free model obtaining less competitive results on classic counting metrics, can instead defeat older learning-based CAC models on the proposed metrics.
@@
From the mosaic test, it is also interesting to understand how the various models drift the estimate about the positive class $c^{\text{pos}}_{ij}$ when changing the negative image $I_{j}$. In other words, we would like to estimate the confusion of the model when it is requested to count a given class, but objects of another class are present in the image. In Figure 4, we report, for each model, the distribution of the quantity $\Bigl{\{}\frac{|c^{\text{pos}}_{ij}-c_{ii}|}{c_{ii}}\Bigl{\}}_{i,j=1}^{N}$, which encodes the normalized absolute error between the estimate of the model when only presented with the correct class as estimated in the negative test $(c_{ii})$ and the estimate of the model when presented with all the mosaics constructed with positive image $I_{i}$ on the top ($c^{\text{pos}}_{ij}$). As we can notice, although DAVE achieves the lowest drift, it also presents many possibly problematic outliers. This may be due to the verification module, which incorrectly assigns the positive label to the cluster obtained from the objects from the negative image. Despite having a higher mean drift, the other methods – especially CounTX and CLIP-Count – show fewer outliers. This signifies that, although these methods wrongly count the negative classes, the count relative to the positive part of the mosaic is less dependent on the number of negative instances from the bottom image of the mosaic. This insight further raises the attention to effective two-stage methods that, despite being very effective in discriminating the different classes, may catastrophically fail in some specific scenarios.
@@
In Figure 5, we show some qualitative results from the various methods examined through the lenses of the mosaic test. The lack of a proper understanding of the correct class to count is largely deducible from the reported density maps. While many methods correctly estimate the positive class on the top of the mosaics, most of the models give a non-zero count prediction of the negative instances from the negative images. We can notice how DAVE, in the first and last rows, correctly ignores the negative instances, exactly estimating zero objects. However, its second-stage verification approach may occasionally fail in a catastrophic manner, assigning the label to all the instances of the negative class, as shown in the second reported example. Although the mosaic test already gives many insights into the failure modes of the probed SOTA models, in the supplementary material, we also report some qualitative results from the negative test.
@@ 5 Conclusions
In this paper, we introduced Prompt-Aware Counting (PrACo), a novel benchmark composed of two targeted test suites specifically designed to address the limitations of current evaluation systems for prompt-based class-agnostic counting (CAC).
Our evaluation of several recent state-of-the-art CAC methods revealed significant performance gaps, particularly in the models’ ability to handle negative and multi-class scenarios. While two-stage models like DAVE generally outperformed one-stage models on the negative test, they still exhibited notable weaknesses, particularly in the mosaic test, where they sometimes struggled with false positives from the negative images.
@@ Appendix A Derivation of Counting Precision and Recall
In this section, we better detail the derivation of Eqs. 4 and 5 within the paper, i.e., the formulas for computing counting precision and counting recall starting from the inferred quantities $c^{\text{pos}}$ and $c^{\text{neg}}$ concerning the mosaic test. For ease of notation, here we omit the indexes $i,j$ from all the involved quantities, given that we focus on a single mosaic.
As stated in the paper, we deal with counting rather than detection. Therefore, we do not know the exact nature of each inferred instance, i.e., we cannot assign a correct/incorrect label to each different detected object. We can anyway directly estimate the total amount of true positives (TPs), false positives (FPs) and false negatives (FNs) by employing the counting model outputs. We perform the following assumption:
  • For the positive image (the top one in the mosaic),
@@
    where $\tilde{c}$ is the ground truth of the positive class. In fact, if the model predicts fewer objects than the ground truth, all the predicted objects are considered correct, and the remaining ones are FNs. Conversely, if the model predicts more objects than the ground truth, only $\tilde{c}$ objects are correct, and the remaining contribute to the FPs. This situation for FNs and FPs can be directly derived from Eq. 7. Indeed, given that $c^{\text{pos}}=\text{FP}^{\text{pos}}+\text{TP}^{\text{pos}}$, it follows that
@@
  • For the negative image (the bottom one in the mosaic), the situation is simpler, given that all the contributions inferred by the model are FPs, as the TPs are identically zero, and thus also the FNs:
@@
With these quantities defined, we can introduce counting precision and counting recall, starting from their definitions in terms of TPs, FPs, and FNs.
@@ A.1 Counting Precision
We start from the definition of precision, which is the following:
@@
Considering that the TPs and FPs are the sums of the respective contributions from the positive and negative parts of the mosaic – i.e. $\text{TP}=\text{TP}^{\text{pos}}+\text{TP}^{\text{neg}}$ and $\text{FP}=\text{FP}^{\text{pos}}+\text{FP}^{\text{neg}}$ – we obtain the precision expressed in terms of the quantities computed in Eqs. 7, 8, 10, and 11. Substituting and simplifying, we obtain
@@
which we can rewrite in simpler manner as
@@ A.2 Counting Recall
The same idea used for deriving the counting precision can also be employed to compute the counting recall. The recall is defined as
@@
Even in this case, TPs and FNs are the sums of the respective contributions from the positive and negative parts of the mosaic – i.e. $\text{TP}=\text{TP}^{\text{pos}}+\text{TP}^{\text{neg}}$ and $\text{FN}=\text{FN}^{\text{pos}}+\text{FN}^{\text{neg}}$. We obtain the precision expressed in terms of quantities computed in Eqs. 7, 9, 10, and 12. Substituting and simplifying, we obtain
@@ Appendix B Derivation of Normalized Mean of Negative predictions (NMN)
NMN, as reported in the paper, is the absolute counting error computed by prompting the model with the negative classes normalized by the ground truth of the positive class. Formally, the main involved quantity computed for each image $I_{i}$ prompted with the negative class $P_{j}$ is given by
@@ Appendix C DAVE Inference Details
In Table 3, we report the model’s behavior with and without modification to the inference strategy and the original and changed $\tau$ parameter. As we can notice, the clustering threshold does not affect the negative test, where only one object cluster is always found. Our modification, which injects positive classes as a reference, originates a strong model from the negative test perspective, with an NMN of only 0.08. Concerning the mosaic model, the lowering of the $tau$ threshold, together with the improved inference procedure, helps raise the counting precision and, in turn, the F1 by more than 12% with respect to the original implementation.

### v2

@@ 3.3 Mosaic test
The mosaic test assesses images containing multiple object categories by providing the model with textual prompts referring only to the positive class, emulating very common real-case scenarios. Thus, the object instances belonging to the negative classes serve as distractors to the positive class. To overcome the lack of suitable annotated CAC datasets having multi-class images, we artificially build a new set of such images. Specifically, starting from existing single-class CAC datasets, we create mosaicked images by stitching together pairs of positive and negative single-class images, similarly to [2].
Formally, we consider all the possible pairs of images $\{I_{i},I_{j}\}_{\begin{subarray}{c}i,j=1\\ j\neq i\end{subarray}}^{N}$, and the positive class $P_{i}$ of the image $I_{i}$. We call $I_{i}$ and $I_{j}$ the positive and the negative images, respectively. We create new mosaicked images $I_{ij}^{\text{mosaic}}=\text{vstack}(I_{i},I_{j})$ by vertically stitching together the positive image, placed at the top, with the negative one, placed at the bottom. Then, the CAC model $\mathcal{M}$ is tasked to predict the following quantities concerning the positive class $P_{i}$:
$c_{ij}^{\text{pos}},c_{ij}^{\text{neg}}=\mathcal{M}(I_{ij}^{\text{mosaic}},P_{i}),$ (4)
@@
where $c_{ij}^{\text{pos}}$ and $c_{ij}^{\text{neg}}$ are the positive and negative class counts relative to the top and the bottom parts of the mosaic, respectively. This can be easily obtained from most density-based models, as it is sufficient to cut the output density map in half along the y-axis and integrate them separately to obtain the respective positive and negative counts. It is worth noting that negative images influence not only the negative class count $c_{ij}^{\text{neg}}$, but also the positive class count $c_{ij}^{\text{pos}}$. Indeed, the goal of this test is to evaluate the robustness of the model to count the positive class described by the prompt in the presence of distractors, i.e., objects from different categories within the same image. Similarly to the negative-label test, we ideally want $c_{ij}^{\text{pos}}$ to converge to the ground truth count, while $c_{ij}^{\text{neg}}$ should ideally approach zero. This means the model (i) should not be distracted by negative examples when attending to positive instances, and (ii) should refrain from counting any negative instances in the mosaic. We summarize this inference procedure in the right part of Fig. 2.
@@
Since each mosaic includes both positive and negative instances, this test bears similarities to object detection evaluation, where both correct and incorrect detections can occur, and performance is typically assessed using precision, which reflects the proportion of correct predictions among all model outputs, and recall, which represents the percentage of correct instances retrieved from the total ground-truth. However, differently from object detection, in our scenario, we do not have clear indications about the correctness of each proposed instance, given that we do not know – and we are not requested to know – precise instance localization within the image, as we are only interested in the final aggregated count for the provided object class. Such information could be partially recovered from the two output counts $c_{ij}^{\text{pos}}$ and $c_{ij}^{\text{neg}}$ by making some reasonable assumptions that we present in the following, where we introduce precision-recall metrics shaped for prompt-based CAC.
Counting Precision (CntP) and Counting Recall (CntR).
@@
Under the above considerations and assumptions, we obtain that counting precision (CntP) and counting recall (CntR) can be defined as follows:
@@
We refer the reader to the supplementary material for the derivation of such expressions from the idea in Fig. 3.
@@
Counting Balanced F1-score (CntF1).
A common way to aggregate precision and recall to obtain a single indicator is the F1-score, which is defined as the harmonic mean of the two metrics. We therefore derive a counting F1-score (CntF1) indicator, as follows:
@@ 4.1 Dataset and Methods
Our benchmark relies on FSC-147 [25], a widely used dataset for CAC containing 6,135 images with objects belonging to 147 classes. Labels include dots over the object centroids, bounding boxes for three object exemplars, and object category textual names (see also Sec. 2). Although most images contain objects belonging to a single category, we filter out the few multi-class images by following the approach in [23]. This prevents interferences with our proposed tests, ensuring that no false positives arise from other object classes present in the same image.
@@
We employed the original code and pre-trained models provided by the authors, maintaining their image pre-processing pipelines, hyperparameters, and prompting schema. We only needed to adjust the DAVE inference procedure to behave correctly on our benchmark. In the supplementary material, we provide further details on these changes along with the PrACo performance of the original DAVE implementation.
@@
Figure 5: This figure shows, for each model, the output density maps for three different (mosaic, input prompt) pairs. The count reported in the blue box is $c^{\text{pos}}_{ij}$, while the count reported in the dark orange box corresponds to $c^{\text{neg}}_{ij}$. We can notice how the models often misidentify instances from the negative image in the mosaic, though most accurately estimate the positive instances in the upper part.
@@ 4.2 Results
We report the results for both the test and validation splits in Tab. 1 and Tab. 2, respectively. As we can notice, although the methods achieve remarkable performance on standard counting error measures (MAE and RMSE), the outcome is quite heterogeneous on the PrACo metrics. Notably, the negative test on one-stage methods like CounTX, CLIP-Count, and VLCounter shows that the average negative count is comparable to the ground-truth count of the correct class ($\text{NMN}\approx 1$), with CLIP-Count and VLCount achieving $\text{NMN}>1$. This trend is further confirmed by the PCCN metric, where CLIP-Count has the correct estimate nearer to the ground-truth only 38% of the times on the test set. Instead, the two-stage detectors TFPOC and DAVE achieve the best performance on the negative test. The same trends are mostly confirmed on the mosaic test. Specifically, while the CntR metric only increments by around 12% between the worst-performing method (CounTX) and the best-performing one (DAVE), the CntP shows a notable increase, from 0.517 of VLCounter to 0.843 of DAVE, which is an improvement of around 63%. This shows how the methods, despite being mostly able to count the correct class, have a highly heterogeneous behavior with respect to the negative images within the mosaics. It is also worth noticing that more recent methods do not always improve on the PrACo metrics, meaning that this aspect is still largely underestimated. It is interesting to observe how TFPOC, a training-free model obtaining less competitive results on classic counting metrics, can instead defeat older learning-based CAC models on the proposed metrics.
@@
From the mosaic test, it is also interesting to understand how the various models drift the estimate about the positive class $c^{\text{pos}}_{ij}$ when changing the negative image $I_{j}$. In other words, we would like to estimate the confusion of the model when it is requested to count a given class, but objects of another class are present in the image. In Fig. 4, we report, for each model, the distribution of the quantity $\Bigl{\{}\frac{|c^{\text{pos}}_{ij}-c_{ii}|}{c_{ii}}\Bigl{\}}_{i,j=1}^{N}$, which encodes the normalized absolute error between the estimate of the model when only presented with the correct class as estimated in the negative test $(c_{ii})$ and the estimate of the model when presented with all the mosaics constructed with positive image $I_{i}$ on the top ($c^{\text{pos}}_{ij}$). As we can notice, although DAVE achieves the lowest drift, it also presents many possibly problematic outliers. This may be due to the verification module, which incorrectly assigns the positive label to the cluster obtained from the objects from the negative image. Despite having a higher mean drift, the other methods – especially CounTX and CLIP-Count – show fewer outliers. This signifies that, although these methods wrongly count the negative classes, the count relative to the positive part of the mosaic is less dependent on the number of negative instances from the bottom image of the mosaic. This insight further raises the attention to effective two-stage methods that, despite being very effective in discriminating the different classes, may catastrophically fail in some specific scenarios.
@@
In Fig. 5, we show some qualitative results from the various methods examined through the lenses of the mosaic test. The lack of a proper understanding of the correct class to count is largely deducible from the reported density maps. While many methods correctly estimate the positive class on the top of the mosaics, most of the models give a non-zero count prediction of the negative instances from the negative images. We can notice how DAVE, in the first and last rows, correctly ignores the negative instances, exactly estimating zero objects. However, its second-stage verification approach may occasionally fail in a catastrophic manner, assigning the label to all the instances of the negative class, as shown in the second reported example. Although the mosaic test already gives many insights into the failure modes of the probed SOTA models, in the supplementary material, we also report some qualitative results from the negative test.
@@ 5 Conclusions
In this paper, we introduced Prompt-Aware Counting (PrACo), a novel benchmark composed of two targeted test suites specifically designed to address the limitations of current evaluation systems for prompt-based CAC.
Our evaluation of several recent state-of-the-art CAC methods revealed significant performance gaps, particularly in the ability of the models to handle negative and multi-class scenarios. While two-stage models like DAVE generally outperformed one-stage models on the negative test, they still exhibited notable weaknesses, particularly in the mosaic test, where they sometimes struggled with false positives from the negative images.
@@
Acknowledgments
This work was partially funded by: Spoke 8, Tuscany Health Ecosystem (THE) Project (CUP B83C22003930001), funded by the National Recovery and Resilience Plan (NRRP), within the NextGeneration Europe (NGEU) Program; SUN – Social and hUman ceNtered XR (EC, Horizon Europe No. 101092612); PNRR-M4C2 (PE00000013) ”FAIR-Future Artificial Intelligence Research” - Spoke 1 ”Human-centered AI”, funded under the NextGeneration EU program; Italian Ministry of Education and Research (MUR) in the framework of the FoReLab projects (Departments of Excellence).
@@ Appendix A Derivation of Counting Precision and Recall
In this section, we provide a more detailed explanation of the derivation of Eqs. 4 and 5 from the paper, specifically the formulas for calculating counting precision and counting recall based on the inferred quantities $c^{\text{pos}}$ and $c^{\text{neg}}$ in the context of the mosaic test. To simplify the notation, we omit the indices $i,j$ from all the involved quantities, as our focus is on a single mosaic.
As stated in the paper, we deal with counting rather than detection. Therefore, we do not know the exact nature of each inferred instance, i.e., we cannot assign a correct/incorrect label to each different detected object. However, we can still estimate the total number of true positives (TPs), false positives (FPs), and false negatives (FNs) directly from the outputs of the counting model. We make the following assumptions:
  • For the positive image (the top part of the mosaic),
@@
    where $\tilde{c}$ is the ground truth of the positive class. Indeed, if the model predicts fewer objects than the ground truth, all the predicted objects are considered correct, and the remaining ones are FNs. Conversely, if the model predicts more objects than the ground truth, only $\tilde{c}$ objects are correct, and the remaining contribute to the FPs. This situation for FNs and FPs can be directly derived from Eq. 8. In fact, given that $c^{\text{pos}}=\text{FP}^{\text{pos}}+\text{TP}^{\text{pos}}$, it follows that
@@
  • For the negative image (the bottom part of the mosaic), the situation is simpler, given that all the contributions inferred by the model are FPs, as the TPs are identically zero, and thus also the FNs:
@@
With these quantities defined, we can introduce the counting precision and the counting recall, starting from their definitions in terms of TPs, FPs, and FNs.
@@ A.1 Counting Precision
We start with the definition of precision, which is the following:
@@
Considering that the TPs and FPs are the sums of the respective contributions from the positive and negative parts of the mosaic – i.e., $\text{TP}=\text{TP}^{\text{pos}}+\text{TP}^{\text{neg}}$ and $\text{FP}=\text{FP}^{\text{pos}}+\text{FP}^{\text{neg}}$ – we obtain the precision expressed in terms of the quantities computed in Eqs. 8, 9, 10 and 12. Substituting and simplifying, we obtain:
@@
which we can rewrite in a simpler manner as:
@@ A.2 Counting Recall
The same idea used for deriving the counting precision can also be employed to compute the counting recall. The recall is defined as:
@@
Even in this case, TPs and FNs are the sums of the respective contributions from the positive and negative parts of the mosaic – i.e., $\text{TP}=\text{TP}^{\text{pos}}+\text{TP}^{\text{neg}}$ and $\text{FN}=\text{FN}^{\text{pos}}+\text{FN}^{\text{neg}}$. We obtain the precision expressed in terms of quantities computed in Eqs. 8, 10, 11 and 13. Substituting and simplifying, we obtain:
@@ Appendix B Derivation of Normalized Mean of Negative predictions (NMN)
NMN, as reported in the paper, is the absolute counting error computed by prompting the model with the negative classes normalized by the ground truth of the positive class. Formally, the main involved quantity computed for each image $I_{i}$ prompted with the negative class $P_{j}$ is given by:
@@ Appendix C DAVE Inference Details
In Tab. 3, we report an ablation study about the model’s behavior (i) with and without modification to the inference strategy, and (ii) the original and changed $\tau$ parameter. As we can notice, the clustering threshold does not affect the negative test, where only one object cluster is always found. Our modification, which injects positive classes as a reference, originates a strong model from the negative test perspective, with an NMN of only 0.08. Concerning the mosaic model, the lowering of the $tau$ threshold, together with the improved inference procedure, helps raise the counting precision and, in turn, the counting F1-score by more than 12% with respect to the original implementation.